OpenStudy (anonymous):

What are the first five terms of the sequence given by the formula an=4n+1

OpenStudy (anonymous):

What is n initialized to?

OpenStudy (anonymous):

put \(n=-1\) get \(4\times 1+1=5\) then put \(n=2\) and so on

OpenStudy (anonymous):

Take the formula an=4n+1 Plug in the subscript n as 1,2,3,4 to find the sequence a(1)=4(1)+1= 5 a(2)=4(2)+1= 9 a(3)=4(3)+1= 13 so on...

OpenStudy (anonymous):

Understand, any q's?

OpenStudy (anonymous):

So then, your first term is 5,9,13, ... Get it :)

OpenStudy (anonymous):

@STEWARTEMILY ??